[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jun 28, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jun 28 17:58:53 UTC 2017


HSU Shift Log for Wed Jun 28 2017 13:58:46 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: GOES-16 not available in NAWIPS or AWIPS; data is  
available on other satellite web pages such as NASA MSFC web page.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170628/400b7b95/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list